OpenStudy (anonymous):

The tangent line to the graph of a function g at the point x = -7 is y = -(3x+25). Find the tangent line to the graph of y = x g(x) at x = -7. Put your answer into slope-intercept form. I can't seem to find g(7) since I need to use the product rule.

OpenStudy (aum):

\[ y = x g(x) \\ \frac{dy}{dx} = xg'(x) + g(x) \\ \left[\frac{dy}{dx}\right]_{x=-7} = (-7)g'(-7) + g(-7) \]The equation of the tangent to g(x) at x = -7 is: y = -(3x+25) The slope of the tangent is -3 and therefore \(\large g'(-7) = -3\) When x = -7, y = -(3*(-7) + 25) = -(-21 + 25) = -4. Thus, (-7, -4) is a point on the tangent AND g(x). \(\large g(-7) = -4\) \[ \left[\frac{dy}{dx}\right]_{x=-7} = (-7)g'(-7) + g(-7) = (-7)(-3)+(-4) = 21-4=17\\ y = x g(x) \\ \text{ } \\ \left[y\right]_{x=-7} =(-7)g(-7) = (-7)(-4) = 28 \\ \text{ } \\ \text{Equation of tangent is: } y = mx + b \\ y = 17x + b \\ \text{When } x = -7, y = 28 \\ 28 = 17(-7) + b \\ 28 = -119 + b \\ b = 28+119 = 147 \\ y = 17x + 147 \]